OpenStudy (anonymous):

Help please!!!! 2.) What is the solution to the proportion below? 8/9=m/72

jimthompson5910 (jim_thompson5910):

Hint: cross multiply to get 8/9=m/72 8*72 = 9m

OpenStudy (anonymous):

thank you!!:)

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

8/9=m/72 8*72 = 9m 576/9=m m=64

jimthompson5910 (jim_thompson5910):

perfect

OpenStudy (anonymous):

:)